A factory manufacturing light-emitting diode (LED) bulb claims that their light bulbs last for 50,000 hours on the average. To confirm if this claim was valid, a quality control manager got a sample of 50 LED bulbs and obtained a mean lifespan of 40,000 hours. The standard deviation of the manufacturing process is 1000 hours. Do you think that the claim of the manufacturer is valid? Use the 95% confidence level.


1
Expert's answer
2021-02-02T05:24:20-0500

The provided sample mean is "\bar{x}=40000" and the sample standard deviation is "s=1000."

The sample size is "n=50." The number of degrees of freedom are "df=n-1=49," and the significance level is "\alpha=0.05."

The following null and alternative hypotheses need to be tested:

"H_0: \mu=50000"

"H_1: \mu\not =50000"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the critical value for a two-tailed test is "t_c=2.009575."

The rejection region for this two-tailed test is "R=\{t:|t|>2.009575\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{40000-50000}{1000/\sqrt{50}}=-70.711"

Since it is observed that "|t|=70.711>2.009575=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"  is different than 50000, at the 0.05 significance level.

Using the P-value approach: The p-value "df=49, t=-70.711," "two-tailed," is "p=0," and since "p=0<0.05," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"  is different than 50000, at the 0.05 significance level.